A metal fabricator produces connecting rods with an outer diameter that has a 1 ± 0.02 inch specification. A machine operator takes several sample measurements over time and determines the sample mean outer diameter to be 1.004 inches with a standard deviation of 0.008 inch.

Calculate the process capability index for this example. (Round your answer to 3 decimal places.)

Process capability index            ???
Answer #1

ANSWER:

Process capability index = 0.667

CALCULATION:

We have,

Mean, M = 1.004 inch
Upper specification limit, USL = 1 + 0.02 = 1.02 inch
Lower specification limit, LSL = 1 - 0.02 = 0.98 inch
Standard Deviation, = 0.008 inch

The formula to work out the process capability index is as below

= Min{(1.004-0.98)/(3*0.008), (1.02-1.004)/(3*0.008)}

= Min (1, 0.67)

= 0.667

Therefore, the process capability index is 0.667
Answer #2

usht upper specification limit = 1+0-02 = 1.02 h5h + hower specification limit - 1-0-02-0-98 Mean (cle) = 1.004 Standard devi